OpenStudy (anonymous):

Part 1: Find the area of a rectangular object which has a length of 2 radical 3 inches and a width of 3 radical 6 inches. (3 points) Part 2: Explain, in complete sentences, how you arrived at the simplified answer and give the final solution in simplified radical form. (2 points) Part 3: What type of object in your home or school might this be? (1 point)

OpenStudy (anonymous):

A = 2√3 * 3√6 Multiply rational numbers together, and radicals together = (2*3) * (√3√6) = 6 * √18 Simplify radical by finding any square factors inside radical = 6 * √(9*2) = 6 * √9√2 = 6 * 3√2 = 18√2 http://answers.yahoo.com/question/index?qid=20110612125542AAIYNhc